Ask
your own question, for FREE!
Probability
14 Online
Pls help:) A box contains 2 red cards, 3 blue cards and 4 yellow cards. Mary repeats drawing one card at a time randomly from the box with replacement until a red card is drawn. Find the probability that Mary needs at least three draws.
Still Need Help?
Join the QuestionCove community and study together with friends!
@kryton1212
thanks
The probability of drawing a red on any draw is \(p= 2/9.\) The probability of successfully drawing a red at the kth draw is $$ P(X=k)=(1-p)^{k-1}p $$ The probability that at least 3 draws is required is $$ P(X \ge 3)=1-P(X<3)=1-P(X\le2)=1-\left[ (p+(1-p)p\right ] $$ Here we used the Geometric Distribution: http://en.wikipedia.org/wiki/Geometric_distribution
Is this the full question?
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
6 hours ago
0 Replies
0 Medals
3 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
1 day ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal